# “Whose Side Are You On?” Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection

### v2

@@ Abstract
Existing approaches to classifying ideology of online content are limited in that they require extensive human effort, labeling of large datasets, and are not able to adapt to evolving ideological contexts. This paper explores the potential of Large Language Models (LLMs) for classifying the political ideology of online content in the context of the two-party US political spectrum through in-context learning (ICL). Our extensive experiments involving demonstration selection in label-balanced fashion, conducted on three datasets comprising news articles and YouTube videos, reveal that our approach significantly outperforms zero-shot and traditional supervised methods. Additionally, we evaluate the influence of metadata (e.g., content source and descriptions) on ideological classification and discuss its implications. Finally, we show how providing the source and description influences the LLM’s classification.
@@ 1 Introduction
Social media platforms are one of the leading sources of information for the majority of Americans Barthel et al. (2020). Observers worry that content on these platforms can cause political extremism Tufekci (2018); Roose (2019), and facilitate the formation of ideological filter bubbles that restrict exposure to diverse viewpoints Pariser (2011). Empirical evidence for these worries is inconsistent Haroon et al. (2023); Hosseinmardi et al. (2021); Chen et al. (2023); Nyhan et al. (2023) in part because scholars use different methodologies (audits versus observational data) and in part because of a lack of robust (i.e., highly accurate across domains) and scalable (i.e., requiring minimal human supervision) methods for estimating political ideology of the content users consume.
@@
In this paper, we advance on literature showing that Large Language Models (LLMs) can be applied to ideology estimation of text in the context of the two-party US political spectrum (democrat/liberal and republican/conservative). They can also adapt to new and evolving information by simple adjustments of the provided prompt. Thus, LLMs can serve as a compelling choice for robust and scalable ideology estimation across diverse text domains and platforms.
@@
Past applications of LLMs in ideology estimation Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023) have primarily focused on zero-shot approaches, utilizing pre-trained LLMs without employing more nuanced few-shot in-context learning (ICL) strategies. Our work advances these prior studies by proposing a modified version of coverage-based ICL that selects demonstrations in a class-balanced manner, significantly improving performance over simple zero-shot prompting and coverage-based ICL. We expand on this by conducting an extensive study through rigorous experiments on three distinct datasets on the capabilities of various LLMs (Llama2-13B, Mistral-7B, and GPT-4o) in estimating the ideology of text. Furthermore, we analyze the effect additional metadata such as content source and description has on the classifications of the LLM. Our analyses provides critical insights into the limitations and potentials of LLMs for nuanced ideological analysis.
@@ 2 Related Work
Prior approaches to ideology estimation have generally fallen under three types. (1) Manual annotations by experts Otero (2019) which, although accurate and robust, are not scalable due to the ever-increasing volume of social media content. (2) Training supervised classifiers Baly et al. (2020) which, although highly scalable, is limited by a need for large volumes of human-labeled data and have difficulty adapting to dynamic and ever evolving ideological contexts (e.g., cross-platform estimation, emerging news events, changes in the political environment). (3) Utilizing social media which uses data from external platforms such as Twitter Le et al. (2017); Barberá (2015); Wojcieszak et al. (2023); Eady et al. (2020) or Reddit Lai et al. (2024) to assess how a given piece of content is shared among users of different ideologies to estimate the ideology of the content itself. This approach is limited in that it requires the content to have been shared by a representative number of users and is agnostic of the nature of the content itself. These limitations prevent these prior approaches from being robust and scalable. On the other hand, in contrast with prior work that uses LLMs for ideology classification mostly in a zero-shot manner Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023), we introduce a class-balanced, coverage-based ICL framework that achieves significant performance improvements over the zero-shot setting.
3 Proposed Approach
@@
Our research problem can be defined as follows: Given content that has an underlying ideological leaning (e.g., a YouTube video, a news article), we aim to determine if the content is liberal, neutral, or conservative. Next, we describe our proposed approach for LLM based ideology classification.
@@
Algorithm 1 presents our proposed ICL solution framework. Given a LLM , a training set of content with ideology class labels (Liberal, Neutral, Conservative) , and number of demonstrations , our task is to classify a given query (an online post, video, or news article) as either liberal, neutral, or conservative. Our approach first builds a smaller candidate set of potential demonstrations and computes a relevance score for each demonstration in the candidate set. From the sorted candidate set, we select demonstrations equally from across all three ideology classes, i.e., only demonstrations can be selected from each ideology class. Thus, we skip over demonstrations with higher scores if demonstrations have already been selected from the class they belong to. The instruction prompts used can be found in Appendix A.
3.1 Test Query Specifications
We specify the test query to be classified as an online post, video, or news article that has a title, source, and a description. An LLM can leverage information from all three components to aid in classification, however, limitations such as context-window sizes and longer compute times for longer text make it worthwhile to explore the minimum amount of information that provides the best accuracy. As such, our experiments rely on four combinations: title; title and source; title and description; and title, source, and description.
3.2 Demonstration Selection
We select demonstrations for in-context learning (ICL) to augment LLM task performance. We utilize Set-Cover Vazirani (2003) to maximize coverage across the candidate pool, where BertScore-Recall (BSR) is chosen as the coverage metric. This strategy is known as Set-BSR, and has been shown to be successful for ICL with multiple classes (Gupta et al., 2023). We also ensure there is an equal number of demonstrations selected from each ideology. To validate this design choice, a comparison of random and balanced demonstration selection can be found in Appendix D.
@@ 3.2 Demonstration Selection
4 Experimental Setup
4.1 Datasets
@@
For our evaluation, we rely on three distinct datasets, which have been used in prior work on ideology classification.
@@
YouTube Slant. We use the 49,894 political videos from the YouTube dataset by Haroon et al. (2023) which estimates the ideology of videos using a Twitter-based slant estimation approach Barberá (2015) on a scale of -1 (most liberal) to +1 (most conservative). We label videos with slant less than -0.33 as Liberal, greater than +0.33 as Conservative, and in-between as Neutral. For the ICL demonstration pool, we uniformly randomly sample 20,000 candidates from the full test set. The test set consists of 1,000 videos sampled as to maintain an equal balance across ideologies in testing.
@@
Ad Fontes News. This dataset consists of US news articles that have been labeled by expert human annotators on Ad Fontes Media Bias Panel Otero (2019). Articles with scores lower than -14.0 are labeled as Liberal, greater than +14.0 as Conservative, and in-between as Neutral. The training set consists of 33,396 news articles from which we build a candidate set of 8,000 using the Set-BSR approach and the test set contains 1,000 news articles with a balanced distribution of ideology.
@@
AllSides News. The dataset contains expert human annotations from AllSides111http://www.allsides.com/media-bias/, which relies on several human annotation approaches to rank news articles in the US such as blind surveys, expert and independent reviews. We use the same training and test sets used by Baly et al. (2019). The training set consists of 26,590 news articles, from which we build a candidate set of 8,000 using the Set-BSR approach. The test set consists of 1,300 videos.
While in our main experiments the candidate set size is large, we show in Appendix C that smaller candidate set sizes yield similar results.
4.2 LLMs
We use the following LLMs for our experiments:
OpenAI GPT-4o. We use the pre-trained GPT-4o model by OpenAI Hurst et al. (2024).
@@ 4.2 LLMs
Meta Llama-2-13B. We use HuggingFace to run Meta’s Llama-2 model with 13B Touvron et al. (2023) parameters on a remote server with a RTX A6000 GPU, 256GB memory, and 48 cores.
@@
MistralAI Mistral-7B. We use HuggingFace to run MistralAI’s Mistral model with 7B parameters Jiang et al. (2023) on the same machine.
4.3 Baselines
@@
Multilayer Perceptron (MLP). A fully connected neural network that takes a 384-dimension text embedding as input, a dense hidden layer of 512 neurons with the tanh activation function, and a three output softmax final layer (for three ideologies).
@@
RoBERTa Baly et al. (2020). As in Baly et al. (2020), we use the base RoBERTa transformer model and fine-tune on each of the three full training datasets for 50 epochs. Unlike the LLM which only receives a small number of demonstrations, the RoBERTa model sees the full dataset(s).
5 Results
@@
Baselines. The baseline MLP model performed the worst, achieving an accuracy of , , and for the YouTube, Ad Fontes, and AllSides datasets when only the title was provided. The RoBERTa baseline achieved a score of on the YouTube dataset and on the AllSides News dataset, marked improvements from the MLP model, but fared worse on the Ad Fontes dataset, with an accuracy of . This shows that simpler models are not well-suited to the task of ideological estimation, especially when compared to more advanced models.
@@
Zero-shot performance. Figure 1 shows the performance of the three LLMs and baselines across the three datasets. Focusing on the zero-shot setting, we observe that GPT-4o outperforms both Mistral and Llama2, achieving an accuracy of , , and across the YouTube, Ad Fontes, and AllSides datasets. Mistral performs better than Llama2, achieving a accuracy on the YouTube dataset closer to GPT-4o whereas Llama2’s accuracy is comparably low at . On the Ad Fontes dataset, GPT-4o performs better than Mistral at accuracy compared to Mistral’s . Finally, all LLMs perform poorly on the AllSides dataset with GPT-4o performing the best at . Comparing with the baselines, despite training on the full training set unlike LLMs, RoBERTa performed far worse on the Ad Fontes dataset at . In contrast, the MLP performs the worst in the YouTube and Ad Fontes datasets.
@@
Algorithm 1 Improvements. Providing the set of demonstrations selected using the Set-BSR approach to the prompt , Figure 1 also shows how the accuracy of the models improves by increasing the number of demonstrations using Algorithm 1. The trend of GPT-4o outperforming Llama2 and Mistral persists as the number of demonstrations increases. GPT-4o consistently achieves the highest scores across all datasets, suggesting that it is particularly well-suited for ideological estimation tasks. Mistral and GPT-4o showed a steady increase in performance as the number of demonstrations increases, both seeing a and increase in accuracy going from ( to for GPT-4o, and to for Mistral). Llama2 saw increases in accuracy across all datasets from , but grew worse as . We also undertake McNemar’s test for assessing statistical significance of all results in Appendix F.
@@
Including Source and Description increases accuracy but not always. Augmenting the test instance to include source and description increases the performance of the models across all the datasets in the zero-shot setting. Figure 2 shows how model performance is affected by altering to test different combinations of the title, source, and description. We consider the best performing ablation of the baseline MLP and RoBERTa models as described in Appendix B. We observed that the source alone was a strong indicator of ideology in the two news datasets and less so in the YouTube slant dataset. These two datasets comprise news articles from well-known news sources of whom the LLMs are most likely aware of allowing it to predict ideology based on its knowledge of the source’s ideology. The accuracy is poorer in the YouTube slant dataset where the sources can be niche and less popular YouTube channels of whom the LLM is unaware of and performed better when the title and description were provided. For the YouTube dataset, performance for GPT-4o increased by when the source and description was provided along with the title. Llama2 saw an increase of and Mistral saw an increase of when the title and source were provided. When augmenting the title with the source and description, we see significantly higher increases for Ad Fontes, with the performance of GPT-4o increasing by , Llama2 by , and Mistral by . Similar increases are seen for AllSides, where adding just the source improved accuracy. We present and discuss these ablations in more detail in Appendix B, and discuss edge cases where the ideology of the source can be misleading in Appendix E.

### v3

@@ Abstract
The rapid growth of social media platforms has led to concerns about radicalization, filter bubbles, and content bias. Existing approaches to classifying ideology are limited in that they require extensive human effort, the labeling of large datasets, and are not able to adapt to evolving ideological contexts. This paper explores the potential of Large Language Models (LLMs) for classifying the political ideology of online content through in-context learning (ICL). Our extensive experiments involving demonstration selection in label-balanced fashion, conducted on three datasets comprising news articles and YouTube videos, reveal that our approach outperforms zero-shot and traditional supervised methods. Additionally, we evaluate the influence of metadata (e.g., content source and descriptions) on ideological classification and discuss its implications. Finally, we show how providing the source for political and non-political content influences the LLM’s classification.
@@ 1 Introduction
Social media platforms are one of the leading sources of information for the majority of Americans and have grown in centrality over the past few years Barthel et al. (2020). Observers worry that the content recommended and seen on social media platforms can radicalize users, cause extremism Tufekci (2018); Roose (2019), and facilitate the formation of filter bubbles that restrict exposure to diverse viewpoints Pariser (2011).
@@
However, empirical evidence for these worries is inconsistent Haroon et al. (2023); Hosseinmardi et al. (2021); Chen et al. (2023); Nyhan et al. (2023) in part because scholars use different methodologies and approaches (audits versus observational data) and in part because we lack robust (i.e., highly accurate across domains) and scalable (i.e., requiring minimal human supervision) methods for estimating political ideology of the content users consume.111We recognize the need to also estimate political ideology of the users themselves, which has been done using variety of methods Barberá (2015) and which is outside the scope of this work. Systematic assessments of whether (and the extent to which) social media platforms indeed expose users to ideologically biased or politically radical content are not possible without robust and scalable political ideology estimation methods.
Prior approaches to ideology estimation have utilized manual expert annotations Otero (2019), supervised classification models Baly et al. (2020), and social media sharing patterns Le et al. (2017); Barberá (2015); Lai et al. (2024); Wojcieszak et al. (2023); Eady et al. (2020). Manual annotations are highly accurate but resource-intensive and difficult to scale, especially given the growing volume of online content. Supervised classification models, while offering some automation, face challenges such as a reliance on extensive labeled data and difficulty adapting to dynamic ideological contexts. Social media-based methods estimate ideology by analyzing content sharing patterns on platforms such as Twitter or Reddit, but these approaches are hindered by data access limitations Murtfeldt et al. (2024), reliance on widely shared content, and being agnostic to the content itself. These limitations highlight the need for scalable, adaptable, and robust solutions to ideology estimation.
@@
In this paper, we advance on literature showing that Large Language Models (LLMs) can be applied to ideology estimation of text in the context of the two-party US political spectrum ranging from democrat/liberal to republican/conservative, and can address the aforementioned challenges of scalability and robustness. More specifically, pre-trained LLMs can learn new classification tasks in-context at test-time Hardt and Sun (2024), which enables a few-shot learning paradigm that only requires a few annotated samples. Furthermore, LLMs can handle complex language subtleties necessary for accurate/robust ideology estimation. They can also adapt to new and evolving information by simple adjustments of the provided prompt. Thus, LLMs can serve as a compelling choice for robust and scalable ideology estimation across diverse text domains and platforms.
@@
Past applications of LLMs in ideology estimation Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023) have primarily focused on zero-shot approaches, utilizing pre-trained LLMs without employing more nuanced few-shot in-context learning (ICL) strategies. Our work advances these prior studies by proposing a modified version of coverage-based ICL that selects demonstrations in a class-balanced manner, substantially improving performance over simple zero-shot prompting and coverage-based ICL. Moreover, whereas existing works primarily classify political texts with explicit ideology (e.g., tweets from senators), we demonstrate the effectiveness of our approach across diverse content categories, including political news, political non-news, and non-political news content.
In sum, we make the following contributions:
We conduct an extensive study on the capabilities of differently sized LLMs (Llama2-13B, Mistral-7B, and GPT-4o) in estimating the ideology of text. Through rigorous experimentation on three distinct datasets previously labeled by domain experts, we demonstrate these models’ accuracy, adaptability, and robustness in capturing ideological cues across multiple data. Our findings position LLMs as a robust and scalable solution for effective ideology estimation.
We propose a demonstration selection algorithm that utilizes coverage-based ICL and class-balanced selection to achieve state-of-the-art LLM performance on ideology estimation. We show that our approach improves upon zero-shot prompting used in past work Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023).
We analyze the effect of additional metadata, such as content source and description, on LLM ideology predictions, especially for neutral or non-political content. Our results show that metadata can influence model predictions, at times aligning the predicted ideology with that of the source or introducing bias when source and content ideologies diverge. This analysis provides critical insights into the limitations and potentials of LLMs for nuanced ideological analysis.
@@ 2 Related Work
Prior approaches to ideology estimation have generally fallen under three types (1) manual expert annotations Otero (2019), (2) content analysis through supervised classification models Baly et al. (2020), and (3) social media sharing approaches Le et al. (2017); Barberá (2015); Lai et al. (2024); Wojcieszak et al. (2023); Eady et al. (2020).
Manual annotations by experts. Several efforts have been taken by independent parties to identify the political ideology of online news articles and sources such as Ad Fontes Otero (2019), AllSides222https://allsides.com, Media Bias/Fact Check333https://mediabiasfactcheck.com, etc through a manual annotation process involving experts. Whereas manual annotation is accurate and robust (albeit depending on the number and quality of experts employed), it is highly resource-intensive and not scalable. Expert annotation becomes especially challenging as the volume of social media content increases consistently.
Training supervised classifiers. In turn, supervised ideology classification models can provide some form of automation. The advent of deep learning techniques revolutionized this field by enabling more sophisticated models capable of capturing complex linguistic patterns. Iyyer et al. (2014) introduced deep neural networks, such as LSTMs and CNNs, which outperformed traditional models in detecting ideology from unstructured text data. In a similar vein, Baly et al. (2020) used a BERT model trained on political news articles to show improvements over existing state-of-the-art supervised learning approaches. However, they are limited by a need for large volumes of human-labeled data and have difficulty adapting to dynamic and ever evolving ideological contexts (e.g., cross-platform estimation, emerging news events, changes in the political environment).
Utilizing social media data. Social media sharing ideology estimation approaches utilize data from external platforms such as Twitter Le et al. (2017); Barberá (2015); Wojcieszak et al. (2023); Eady et al. (2020) or Reddit Lai et al. (2024) to assess how a given piece of content is shared among users. By identifying left- and right-leaning elites (e.g., politicians, news media) those users follow, those approaches first estimate the ideology of the users, and then estimate the ideology of the shared content by aggregating the ideology of the sharing users. This approach is limited in several ways. First, as it relies on third-party platforms (such as Twitter/X), changes to APIs and data acquisition systems can render this approach untenable. For instance, since Twitter transitioned to X in 2023, it has become challenging for researchers to collect sharing data at scale without incurring significant costs Murtfeldt et al. (2024); Poudel and Weninger (2024). Second, to accurately estimate the ideology of content (e.g., a YouTube video), the content needs to be shared extensively, which might not happen for niche videos with small viewership. Finally, this approach is agnostic to the nature of the content itself and only looks at how it is shared by users. Owing to these limitations, these approaches may not be robust and reliable.
Using Large Language Models. LLMs have achieved state-of-the-art performance for several tasks across multiple language domains (e.g., legal transcripts Colombo et al. (2024) and medical data Yang et al. (2022)) and contexts (e.g., identifying misinformation Hoes et al. (2023), hate speech Huang et al. (2023), sentiment Heseltine et al. (2024), or relevance, topics, and frames in tweets and news articles Gilardi et al. (2023).
As aforementioned, prior works have utilized LLMs to classify the ideology of text Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023). As we show in Table 1, our work differs from these in several key ways. First, all these approaches are zero-shot methods that only utilize the pre-trained LLMs as is, and do not employ optimal and efficient few-shot in-context learning (ICL) approaches to improve accuracy and adapt LLMs to multiple domains. We propose a modified version of coverage-based ICL that selects demonstrations in a class-balanced manner, and improves performance over simple zero-shot prompting and coverage-based ICL. Second, these methods focus on classifying texts from sources with clearly distinct ideology (e.g. tweets from liberal or conservative politicians). We focus on content with differential levels of ideological signal to demonstrate that our LLM-based approach can estimate ideology of political news, political content but not news, and non-political news content as well. We also analyze the potential limitations for ideology estimation of each of these content subcategories. Finally, prior works do not undertake extensive benchmarking of performance across multiple LLMs and diverse datasets. We run experiments on three different datasets (two of news articles and one consisting of YouTube videos) and three different LLMs (open-source and closed-source) of different sizes. Our results shed light on which LLMs are better suited for ideology classification and what data/information augments LLM performance, i.e. whether additional video descriptions or channel information in prompts increases/reduces performance.
3 Problem Formulation
@@
Our research problem can be defined as follows: Given content that has an underlying ideological leaning (e.g., a YouTube video, a news article), we aim to determine if the content is liberal, neutral, or conservative. Next, we describe our proposed approach where we utilize set coverage and balanced shot selection for ideology classification.
3.1 Proposed Approach
@@
Algorithm 1 provides our proposed in-context learning (ICL) framework. Given a LLM , a training set of content with ideology class labels (Liberal, Neutral, Conservative) , and number of demonstrations , our task is to classify a given query (an online post, video, or news article) as either liberal, neutral, or conservative. Our approach first builds a smaller candidate set of potential demonstrations and computes a relevance score for each demonstration in the candidate set. From the sorted candidate set, we select demonstrations equally from across all three ideology classes, i.e., only demonstrations can be selected from each ideology class. Thus, we skip over demonstrations with higher scores if demonstrations have already been selected from the class they belong to. The instruction prompts used can be found in Appendix A.
@@
3.2 Test Query Specifications
We specify the query to be classified as an online post, video, or news article that has a title, source, and a description. An LLM can leverage information from all three components to aid in classification, however, limitations such as context-window sizes and longer compute times for longer text make it worthwhile to explore the minimum amount of information that provides the best accuracy. As such, our experiments rely on four combinations: title; title and source; title and description; and title, source, and description.
| Sample Text |
| Trump tweets COVID-19 video so outrageous it was banned from Facebook and Twitter |
| Selected Demonstrations |
| Twitter removes Trump retweet that falsely claimed COVID-19 ‘has a cure’ |
| Reporter to sue Michigan Gov. Whitmer for allegedly not releasing COVID-19 nursing home data |
| AU Health masters COVID-19 therapy but loses $20 million |
| Facebook to label COVID-19 vaccine posts to combat misinformation, help people find where to get the shot |
| Trump Gives Delusional COVID-19 Press Conference Like A Broken Man |
| Florida Scientist Says She Was Fired For Not Manipulating COVID-19 Data |
| Wisconsin pharmacist arrested for sabotaging COVID-19 vaccines |
| Facebook openly conspiring with FBI to punish Trump supporters for wrongthink |
3.3 Demonstration Selection
We select demonstrations for in-context learning (ICL) to augment the LLM’s performance. We utilize Set-Cover Vazirani (2003) to maximize coverage across the candidate pool, where BertScore-Recall (BSR) is chosen as the coverage metric. This strategy is known as Set-BSR, and has been shown to be successful for ICL with multiple classes (Gupta et al., 2023). Additionally, we ensure that an equal number of demonstrations is selected from each ideology class. A comparison of random and balanced shot selection can be found in Appendix G to validate this design choice.
4 Results and Evaluation
4.1 Experimental Setup
4.1.1 Datasets
@@
For our evaluation, we rely on three different datasets, which have been used in prior work on ideology classification.444While we use a conservatively large candidate set size, we show in Appendix F that even a smaller candidate set size yields similar results.
@@
YouTube Slant Dataset. We sample 645,187 videos from the YouTube dataset by Haroon et al. (2023) which estimates the ideology of videos using the Twitter-based slant estimation approach Le et al. (2017); Barberá (2015); Wojcieszak et al. (2023) on a scale of -1 (most liberal) to +1 (most conservative) and was collected between 2021 and 2022. We use a BERT-based political classifier from Askari et al. (2024) to further identify 49,894 political videos to create our YouTube slant dataset. Videos with slant as were labelled as liberal, as conservative, and everything else as neutral. The training set used for selection consists of 48,895 videos from which 20,000 candidates were selected and ranked using the Set-BSR approach. The test set consists of 999 videos sampled as to maintain an equal balance of ideologies. Overall, we have videos from over 8,523 channels across the entire training and testing sets (See Appendix I for where the slant cut-offs lie respective to popular news channels.).
@@
Ad Fontes News Articles. This dataset consists of news articles that have been labeled by expert human annotators on Ad Fontes Media Bias Panel Otero (2019) from “most extreme left” to “most extreme right.” Each article is rated by at least three experts from across the political spectrum on a scale of -42 to +42 corresponding to extremely left and extremely right news articles, respectively. Additionally, Ad Fontes labels consider articles with scores lower than -14.0 to skew left and articles with scores greater than +14.0 as skewing right. Adopting these cut-offs, we label articles with scores lower than -14.0 as Liberal, scores greater than +14.0 as Conservative, and everything in-between as Neutral. The training set consists of 33,396 news articles from which we build a candidate set of 8,000 using the Set-BSR approach. The test set contains 999 news articles with a balanced distribution of ideology.
@@
AllSides News Articles. The dataset contains expert human annotations from AllSides555http://www.allsides.com/media-bias/media-bias-rating-methods, which relies on several human annotation approaches to rank news articles such as blind surveys, expert reviews (on a scale of -6 to +6), and independent reviews. We use the training and test set of news articles used by Baly et al. (2019) from this dataset. The training set consists of 26,590 news articles, from which we build a candidate set of 8,000 using the Set-BSR approach. The test set consists of 1,300 videos.
4.1.2 LLMs
We use three different publicly available LLMs for the experiments.
OpenAI GPT-4o. We use the pre-trained GPT-4o model by OpenAI Hurst et al. (2024) using a paid API subscription.
@@ 4.1.2 LLMs
Meta Llama-2. We use the Hugging Face library to run Meta’s Llama-2 model with 13B Touvron et al. (2023) parameters on a local machine with a RTX A6000 GPU, 256GB memory, and 48 cores.
@@
MistralAI Mistral. We use the Hugging Face library to run MistralAI’s Mistral model with 7B parameters Jiang et al. (2023) on the same local machine with a RTX A6000 GPU, 256GB memory, and 48 cores.
4.1.3 Baselines
@@
Multilayer Perceptron (MLP). This model is a fully connected neural network that takes a 384-dimension text embedding as input to an initial dense hidden layer of 512 neurons with the tanh activation function, and a three output softmax final layer (for the three ideologies).
@@
RoBERTa. As in Baly et al. (2020), we use the base RoBERTa transformer model and fine-tune on each of the three full training datasets for 50 epochs. Unlike the LLM which only receives a small number of demonstrations, the RoBERTa model sees the full dataset(s).
4.2 Results
@@
Baselines. The baseline MLP model performed the worst, achieving an accuracy of 0.49, 0.38, and 0.41 across the YouTube, Ad Fontes, and AllSides datasets. The RoBERTa baseline achieved a score of 0.65 on the YouTube Dataset and 0.57 on the AllSides News dataset, marked improvements from the MLP model, but fared worse on the Ad Fontes dataset, with accuracy of 0.33. This clearly shows that simpler models are not well-suited to the task of ideological estimation, especially when compared to more advanced transformer-based architectures.
@@
Zero-Shot Learning. Figure 1 shows the performance of the three LLMs and baselines across the three datasets. Focusing on the zero-shot setting, we notice that GPT-4o outperformed both Mistral and Llama2, achieving an accuracy of 0.64, 0.69, and 0.50 across the YouTube, Ad Fontes, and AllSides datasets. Mistral performed better than Llama2, achieving a 0.59 accuracy on the YouTube dataset closer to GPT-4o whereas Llama2’s accuracy was very low at 0.44. On the Ad Fontes dataset, GPT-4o performed 14% better than Mistral at 0.69 accuracy compared to Mistral’s 0.55. Finally, all LLMs performed poorly on the AllSides dataset with GPT-4o performing the best at 0.50, 7% lower than the baseline RoBERTa model. Comparing with the baselines, RoBERTa performed slightly better than GPT-4o at 0.65 and 0.57 on the YouTube Slant and AllSides datasets, respectively. In contrast, the MLP performed the worst in the Ad Fontes and AllSides datasets but fared better than Llama2 in the YouTube slant dataset.
@@
Few-Shot ICL Comparison. Providing the set of demonstrations selected using the Set-BSR approach to the prompt , Figure 1 also shows how the accuracy of the models improved by increasing the number of demonstrations . The trend of GPT-4o outperforming Llama2 and Mistral persisted as the number of demonstrations increased. GPT-4o consistently achieved the highest scores across all datasets, suggesting that it is particularly well-suited for ideological estimation tasks.
Mistral-7B and GPT-4o showed a steady increase in performance as the number of demonstrations increased, both seeing a 5% increase in accuracy going from (0.64 to 0.69 for GPT-4o, and 0.59 to 0.64 for Mistral-7B). Llama2-13B saw marked increases in accuracy across all datasets from , but grew worse as .
@@
Providing Additional Metadata. Augmenting the test instance to include source and description increased the performance of the models across all the datasets in the zero-shot setting. Figure 2 shows how model performance was affected by altering to include the title, the title and source, the title and description, and the title, source, and description. For the YouTube dataset, performance for GPT-4o increased by 7% when the source and description was provided along with the title. Llama2 saw an increase of 5% when the description was provided and Mistral saw an increase of 8% when only the source was provided. When augmenting the title with the source and description, we see higher increases for Ad Fontes, with the performance of GPT-4o increasing by 17%, Llama2 by 21%, and Mistral by 19%. Similar increases are seen for AllSides, where adding just the source pushed the accuracy of the models further. In short, when provided the source information, all three LLMs were able to surpass the baseline RoBERTa model in the zero-shot setting.
Classifying non-political content. Our findings so far have focused on exclusively political content. The improvement in the accuracy of the LLMs when the source is provided suggests that the source carries substantial weight in determining the ideology of content, where the title alone was previously misclassified. This improvement is more evident in the Ad Fontes and AllSides news datasets, where the sources are well-known news organizations, and less so in the YouTube slant dataset, where the sources can be any YouTube channel. To pinpoint the relative role of the title and source in determining the classification, we resample the entire YouTube slant dataset to procure three datasets: 1) political content from well-known news channels (political news), 2) non-political content from well-known news channels (non-political news), and 3) political news from non-news channels, such as influencers or pundits (political non-news). We use the same training set for the YouTube slant dataset and estimate ideology using GPT-4o.
| Political | Political | Non-political | |
| news | non-news | news | |
| Liberal | 314 | 309 | 125 |
| Neutral | 435 | 436 | 207 |
| Conservative | 251 | 255 | 76 |
| Total | 1,000 | 1,000 | 408 |
Breaking this down by the class labels, Figure 3 shows the change in prediction accuracy after adding source information using a heatmap of changes across the three datasets. The heatmap shows which corresponding class was a given label misclassified as before providing the source allowing us to analyze the influence of the source when estimating ideology. Coupling this with qualitative examples from Appendix H provides insight into the influence of the source in classifying ideology.
For political news, we see an improvement in accuracy of content previously misclassified as neutral to liberal (+16%) and conservative (+7%), becoming almost 70% more accurate in both cases. When looking at the kinds of titles re-classified in Table 16 in Appendix H, we see that the titles themselves do not indicate the any ideological leaning. For example, “Haberman says Trump has been diminished but isn’t dead politically” was classified as neutral but when the LLM was informed it was from CNN, it was re-classified as liberal. Similarly, the video titled “Transgender Clinical Psychologist on people de-transitioning | Dr. Erica Anderson” was re-classified as conservative when the LLM learned it came from Newsmax, a conservative outlet, despite the title itself indicating no such political leaning. Cases where the label changed from liberal to conservative or vice-versa showed more nuance. The video titled “Democrats introduce bill to block Trump from holding office” was classified as liberal but then re-classified as conservative when the source Newsmax was provided which was consistent with the content of the video. This is true for the non-political news dataset as well, where the titles do not depict any political inclination as can be seen in Table 17 in Appendix H but the model re-classified it according to the ideology of the news source. Finally, for the political non-news videos, we do not see much movement as the LLM is likely unfamiliar with the sources provided. The cases where we do see movement as seen in Table 18, it is also driven by the source of the video as well. We see some movement towards the neutral as well. For example, the videos titled “The Trouble with the Electoral College” and “Is capitalism actually broken?” are liberal in ideology but when the model learns they are from educative channels CGP Grey and TED-Ed, it re-classifies them as neutral.
4.3 Discussion
Few-shot ICL increases performance. With higher number of shots, we see that accuracy improves across the board for all three LLMs across all three datasets. Except for Llama2-13B which performs best at 4-shots (likely due to its shorter context window length), both GPT-4o and Mistral-7B see continued improvement in accuracy at an average of 5% and 3% across all three datasets.
Supplementing prompt with additional metadata improves performance. The additional metadata consist of the source and description of the content. Just adding the content source increased the accuracy of all the models, suggesting that the models are familiar with the ideological positions of the sources used and leverage that information to improve their ideology estimation. In contrast, the description actually lowers the accuracy when added, likely because it affects the context-window of the input poorly and may not provide more context than the title itself.
Leveraging few-shot ICL and additional metadata for best performance. Combining the few-shot ICL and incorporating the source resulted in the highest gains in GPT-4o’s performance across all three datasets. The model achieved a 10% increase in the YouTube slant dataset (), 18% increase in the Ad Fontes dataset (), and 37% increase in the AllSides dataset (). These improvements demonstrate how GPT-4o leverages details from relevant demonstrations and its existing knowledge of news sources to more accurately reason about the ideology of content.
Multi-modal classification using YouTube thumbnails does not improve GPT-4o’s performance. We experiment with adding video thumbnails as part of the prompt to GPT-4o. Despite the potential of multi-modal classification methods to enhance model performance by incorporating visual information, our experiments reveal that using YouTube thumbnails as an additional modality does not improve GPT-4o’s performance. When comparing the results of GPT-4o with and without the inclusion of YouTube thumbnails, the accuracy of the model dropped to 0.35 from 0.64 on the YouTube slant dataset. This result indicates that, for the problem of ideology estimation from YouTube content, visual information does not contribute additional meaningful signals beyond what is captured in the text and instead can be misleading for the model. However, more extensive analysis and experiments are needed for incorporating multi-modal information for ideology classification and can be undertaken in future work.
Chain-of-thought reasoning is ineffective. We apply chain-of-thought (CoT) reasoning (i.e., prompting the model to think through the task step-by-step) to GPT-4o for our task and see that it doesn’t improve accuracy despite its known benefits Kojima et al. (2024). On the YouTube slant dataset, the model accuracy dropped from 0.64 to 0.62 when CoT reasoning was used. One explanation for this finding is that ideology estimation is primarily reliant on identifying patterns in content that are explicit rather than requiring deep logical steps. The task might not demand the same kind of multi-step reasoning that CoT typically helps with in applications, such as arithmetic or procedural problem-solving.
Accuracy of LLMs is comparable with human annotation. The high accuracy of GPT-4o is comparable to human annotation on the AllSides and Ad Fontes News datasets, both of which are human-annotated. On these, GPT-4o achieved high accuracy, closely aligning with human judgments, exceeding 0.80 in some configurations. For additional validation, a domain expert annotated a sample of 561 videos in the YouTube dataset, finding that GPT-4o agreed with the human annotation 79% of the time. This suggests that GPT-4o is able to effectively mirror human-level ideological classification, even in more informal content. The model’s performance highlights its capacity to generalize patterns and signals used by human annotators, making it a valuable tool for automated ideological estimation.